How long will it take for $1,600 to double if it is invested at 11% annual interest compounded 4 times a year? Enter in exact calculations or round to 3 decimal places.

Answer by mananth(16946)

Principal P = 1600
Amount= 3200
n= number of periods 1
compounded (t) 4
Rate = 11 0.11
n = ln(FV / PV) / ln(1 + r) (r/q))

n=( ln ( 3200 / 1600 )/ln( 1 + 0.03 )
n=( 0.69 )/ ln 1.03
n=( 0.69 / 0.03 )
n= 25.55 quarters
6.39 years